[Hallc_running] [New Logentry] RC daily update

sjjooste at jlab.org sjjooste at jlab.org
Tue Oct 16 13:15:02 EDT 2018

Logentry Text:
Beam has been much better after yesterday's beam studies. They brought beam back
to the hall around 5pm. The accelerator was able to sustain 55uA for more than 10min
without tripping for one of our first runs!
More info here: https://logbooks.jlab.org/entry/3610377

We were down for beam tuning until ~5pm, after which we started the luminosity scan.  For the scan we varied spectrometer settings, targets, and beam currents (between 55u and 5uA in 5 steps). You can find the full run plan for the sweep here: https://hallcweb.jlab.org/wiki/index.php/Run_Plan_for_Kaon_LT_Luminosity_Scan.
We started the scan at 5:30pm and finished around 3am. I would like to thank the swing shift crew (Holly and Rolf) as well as the owl shift crew (Burcu and Hem) for their effort to complete this very involved run plan efficiently.

After the scan the owl shift crew resumed physics data taking. They were able to take two good runs before the accelerator was taken offline for RF recovery.

During RF recovery, we went in controlled access. During this time, Brad and Pavel installed a NDX detector for RadCon in the SHMS hut https://logbooks.jlab.org/entry/3610342. At the same time, Steve Lassiter and Andy Kenyon worked on various issues. They worked on the troublesome HMS dipole LN2 JT-4 valve as planned. Additionally, they used this opportunity to do maintenance work on the HMS and SHMS rotation system. They also investigated the cause behind the loud screeching noise when opening the HMS shutter. Andy traced this issue to a bent rod: not a pressing issue but should be fixed eventually before anything gets damaged.

Around noon we went back to beam permit. As of right now, Hall B, Hall D and the southern linac are still in controlled access.

Current ETA to resume physics running is the start of the swing shift.

As of right now, we have collected 26% of the statistics goal for LH2 on Q2=5.5, W=3.02, theta_Shms=6.49 during 3.25 shifts.
We will resume taking production data on this same kinematic point for another 2-3 shifts, or until we reach 50% of the statistics goal.
We will move to the next kinematic point tomorrow morning.
More info about the Kaon LT status here: https://hallcweb.jlab.org/wiki/index.php/Kaon_LT_Status_Page

After the meeting you can find the talking points here:

To join the Meeting:

To join via Room System:
Video Conferencing System: bjn.vc -or-
Meeting ID : 861439920

To join via phone :
1) Dial:
+1.408.740.7256 (US (San Jose))
+1.888.240.2560 (US Toll Free))
+1.408.317.9253 (US (Primary, San Jose))
(see all numbers - http://bluejeans.com/numbers)

2) Enter Conference ID : 861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3610389
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181016/31f4afba/attachment.html>

More information about the Hallc_running mailing list